# Human-CLAP: Human-perception-based contrastive language–audio pretraining

### v2

@@ 1 Introduction
Various tasks focusing on audio–text relationships, such as text-to-audio (TTA) [liu2023audioldm, 10112585, kreuk2023audiogen], language-queried audio source separation (LASS) [lass, 10752098], and automated audio captioning (AAC) [Drossos_2017_waspaa, shah2023_t6b, greeshma2023_t6a], have been developed. Constructing text representations of audio is a common effective approach across such audio–text tasks, making audio more interpretable for humans [wang2023_t6b]. Representing audio as text descriptions improves the searchability of audio data and allows more people, including those with hearing difficulties, to enjoy audio contents.
CLAP [laionclap2023, elizalde2023clap] has been proposed as a foundation model for connecting audio and text by contrastive learning. By learning the correspondence of audio–text pairs, CLAP creates a shared embedding space between audio and text, enabling wide applications to audio recognition and generation tasks [pmlr-v202-huang23i]. Moreover, CLAP is used not only for these tasks, but also for the evaluation of relevance between audio and text in TTA [liu2024audioldm, majumder2024tango, ghosal2023text, 10731578, ijcai2024p502]. Concretely, the distance between the audio and text embedding vectors obtained by CLAP can be interpreted as a score that quantifies the degree of relevance between the audio and the text, which is called CLAPScore. Although CLAPScore is used for the objective evaluation in TTA tasks [liu2024audioldm, majumder2024tango], the relationship between CLAPScore and human subjective evaluation scores has not been investigated. Considering that TTA aims to generate sounds humans hear, it is necessary to investigate how well the CLAPScore aligns with human subjective evaluation scores. In addition, although CLAP is trained with the assumption that the audio–text pairs in the dataset match exactly, the audio–text pairs also include noisy data, such as text that does not contain all the content in the audio. Therefore, there is a possibility that the distance between audio and text will be close in the embedding space of CLAP, even for inappropriate audio–text pairs. It is also possible to collect cleaner audio–text pairs for model training, but data collection is extremely costly.
@@ 2.1 CLAPScore
CLAPScore is an evaluation metric for the relevance of texts and audio samples in TTA. CLAPScore utilizes CLAP [elizalde2023clap] embeddings and is defined as
@@
where $\textbf{e}^{\textsf{text}}$ and $\textbf{e}^{\textsf{audio}}$ denote the CLAP embedding of text and audio, respectively [liu2024audioldm]. Although CLAPScore is sometimes calculated to take values between $-1$ and $1$, in this paper, we defined negative similarity scores as 00 to compare correlation with human evaluation scores. This definition follows the precedence set by the conventional work on CLIPScore [hessel-etal-2021-clipscore].
@@ 2.2 Dataset of subjective scores for audio–text pairs
We used the RELATE dataset [kanamori_INTERSPEECH2025], which contains subjective evaluation scores of the relevance between audio–text pairs. The RELATE dataset includes natural audio samples from AudioCaps [kim2019audiocaps] and synthesized audio samples generated using AudioLDM [liu2023audioldm], AudioLDM2 [liu2024audioldm], Tango [ghosal2023text], and Tango2 [majumder2024tango].
@@
To select high-quality listeners, we conducted a screening of listeners using anchor samples in the listening test. The RELATE dataset provides three anchor samples, which were mismatched audio–text pairs, in each listening test. We removed listeners who scored higher than 2 on average for anchor samples. This screening method is based on the subjective evaluation conducted by Choi et al. in DCASE 2023 Challenge Task 7 [choi:hal-04178960].
@@ 2.3 Correlation between CLAPScore and Subjective Evaluation Scores
We investigated the relationship between CLAPScore and subjective evaluation score using 2,405 audio–text pairs from test data collected in Section 2.2. Two types of CLAP model were used, models by Wu et al. (LAION CLAP) [laionclap2023] and Elizalde et al. (MS CLAP) [elizalde2023clap]. To evaluate the correlation between CLAPScore and subjective evaluation scores, we used Spearman’s rank correlation coefficient (SRCC), linear correlation coefficient (LCC), and Kendall’s rank correlation coefficient (KTAU).
The CLAPScore calculated with LAION CLAP had an SRCC of $0.280$ and an LCC of $0.294$. Similarly, those calculated with MS CLAP had an SRCC of $0.278$ and an LCC of $0.296$. From these results, CLAPScore has a low correlation with subjective evaluation scores. Such a low correlation is also visualized in the scatter plot of the relationship between CLAPScore and subjective score shown in Fig. LABEL:fig:laion_test. These results indicate that CLAPScore is not necessarily suitable for an objective evaluation metric in TTA.
@@ 3.2 Loss Function
Our Human-CLAP model is trained to minimize a loss function that combines contrastive learning and regression approaches. Regression losses such as the mean square error (MSE) and mean absolute error (MAE) are often used to train models that predict subjective evaluation scores for speech–text pairs or audio–text pairs [tjandra2025aes, baba2024utmosv2, lo19_interspeech]. In this study, we used the regression loss between the target score $a_{i}\in[0,1]$ and the predicted cosine similarity $y_{i}$. We utilize MSE and MAE to calculate such regression losses. MSE is calculated as
@@
We also trained the model with the contrastive learning approach performed in the conventional CLAP model [elizalde2023clap, laionclap2023]. The conventional CLAP is trained by using symmetric cross entropy (SCE) loss [laionclap2023, DBLP:conf/icml/RadfordKHRGASAM21] to construct its audio–text embedding space. We propose the weight SCE (wSCE) loss by adding weights based on the subjective score to the SCE loss:
@@ 4.1 Experimental Setup
We fine-tuned LAION CLAP [laionclap2023] using the subjective evaluation score collected in Section 2. The learning rate was set to $10^{-5}$ and AdamW [loshchilov2018decoupled] was used for optimization, and the batch size was 8. We trained the model for 50 epochs and selected the best model that had the lowest validation loss. The hyperparameters of $L$ (Eq. (6)) were emperically determined to be $[\lambda_{1},\lambda_{2}]=[0.1,1]$ after multiple rounds of hyperparameter tuning. We used RoBERTa [DBLP:journals/corr/abs-1907-11692] and hierarchical token-semantic audio Transformer (HTS-AT) [chen2022hts] for the text and audio encoders, respectively. We split the train data collected in Section 2 into 1,925 and 458 audio–text pairs for model training and validation.
We evaluated the correlation between CLAPScore and subjective evaluation scores using 2,405 audio–text pairs collected from the AudioCaps [kim2019audiocaps] test set in Section 2. We used SRCC and LCC to calculate the correlation between CLAPScore and subjective score, and used MSE to evaluate the error between CLAPScore and subjective score. LAION CLAP [laionclap2023] and MS CLAP [elizalde2023clap] are used as the comparison methods.
@@ 4.2 Results
Model Sub. score $\leq 5$ Sub. score $>5$

### v3

@@ 1 Introduction
Various tasks focusing on audio–text relationships, such as text-to-audio (TTA) [15, 28, 13], language-queried audio source separation (LASS) [17, 21], and automated audio captioning (AAC) [4, 11, 7], have been developed. Constructing text representations of audio is a common effective approach across such audio–text tasks, making audio more interpretable for humans [25]. Representing audio as text descriptions improves the searchability of audio data and allows more people, including those with hearing difficulties, to enjoy audio contents.
CLAP [26, 5] has been proposed as a foundation model for connecting audio and text by contrastive learning. By learning the correspondence of audio–text pairs, CLAP creates a shared embedding space between audio and text, enabling wide applications to audio recognition and generation tasks [9]. Moreover, CLAP is used not only for these tasks, but also for the evaluation of relevance between audio and text in TTA [16, 22, 6, 27, 14]. Concretely, the distance between the audio and text embedding vectors obtained by CLAP can be interpreted as a score that quantifies the degree of relevance between the audio and the text, which is called CLAPScore. Although CLAPScore is used for the objective evaluation in TTA tasks [16, 22], the relationship between CLAPScore and human subjective evaluation scores has not been investigated. Considering that TTA aims to generate sounds humans hear, it is necessary to investigate how well the CLAPScore aligns with human subjective evaluation scores. In addition, although CLAP is trained with the assumption that the audio–text pairs in the dataset match exactly, the audio–text pairs also include noisy data, such as text that does not contain all the content in the audio. Therefore, there is a possibility that the distance between audio and text will be close in the embedding space of CLAP, even for inappropriate audio–text pairs. It is also possible to collect cleaner audio–text pairs for model training, but data collection is extremely costly.
@@ 2.1 CLAPScore
CLAPScore is an evaluation metric for the relevance of texts and audio samples in TTA. CLAPScore utilizes CLAP [5] embeddings and is defined as
@@
where $\textbf{e}^{\textsf{text}}$ and $\textbf{e}^{\textsf{audio}}$ denote the CLAP embedding of text and audio, respectively [16]. Although CLAPScore is sometimes calculated to take values between 1-1 and $1$, in this paper, we defined negative similarity scores as 0 to compare correlation with human evaluation scores. This definition follows the precedence set by the conventional work on CLIPScore [8].
@@ 2.2 Dataset of subjective scores for audio–text pairs
We used the RELATE dataset [10], which contains subjective evaluation scores of the relevance between audio–text pairs. The RELATE dataset includes natural audio samples from AudioCaps [12] and synthesized audio samples generated using AudioLDM [15], AudioLDM2 [16], Tango [6], and Tango2 [22].
@@
To select high-quality listeners, we conducted a screening of listeners using anchor samples in the listening test. The RELATE dataset provides three anchor samples, which were mismatched audio–text pairs, in each listening test. We removed listeners who scored higher than 2 on average for anchor samples. This screening method is based on the subjective evaluation conducted by Choi et al. in DCASE 2023 Challenge Task 7 [3].
@@ 2.3 Correlation between CLAPScore and Subjective Evaluation Scores
We investigated the relationship between CLAPScore and subjective evaluation score using 2,405 audio–text pairs from test data collected in Section 2.2. Two types of CLAP model were used, models by Wu et al. (LAION CLAP) [26] and Elizalde et al. (MS CLAP) [5]. To evaluate the correlation between CLAPScore and subjective evaluation scores, we used Spearman’s rank correlation coefficient (SRCC), linear correlation coefficient (LCC), and Kendall’s rank correlation coefficient (KTAU).
The CLAPScore calculated with LAION CLAP had an SRCC of $0.280$ and an LCC of $0.294$. Similarly, those calculated with MS CLAP had an SRCC of $0.278$ and an LCC of $0.296$. From these results, CLAPScore has a low correlation with subjective evaluation scores. Such a low correlation is also visualized in the scatter plot of the relationship between CLAPScore and subjective score shown in Fig. 1. These results indicate that CLAPScore is not necessarily suitable for an objective evaluation metric in TTA.
@@ 3.2 Loss Function
Our Human-CLAP model is trained to minimize a loss function that combines contrastive learning and regression approaches. Regression losses such as the mean square error (MSE) and mean absolute error (MAE) are often used to train models that predict subjective evaluation scores for speech–text pairs or audio–text pairs [24, 1, 19]. In this study, we used the regression loss between the target score $a_{i}\in[0,1]$ and the predicted cosine similarity $y_{i}$. We utilize MSE and MAE to calculate such regression losses. MSE is calculated as
@@
We also trained the model with the contrastive learning approach performed in the conventional CLAP model [5, 26]. The conventional CLAP is trained by using symmetric cross entropy (SCE) loss [26, 23] to construct its audio–text embedding space. We propose the weight SCE (wSCE) loss by adding weights based on the subjective score to the SCE loss:
@@ 4.1 Experimental Setup
We fine-tuned LAION CLAP [26] using the subjective evaluation score collected in Section 2. The learning rate was set to $10^{-5}$ and AdamW [20] was used for optimization, and the batch size was 8. We trained the model for 50 epochs and selected the best model that had the lowest validation loss. The hyperparameters of $L$ (Eq. (6)) were emperically determined to be $\lambda_{1},\lambda_{2}]=[0.1,1]$ after multiple rounds of hyperparameter tuning. We used RoBERTa [18] and hierarchical token-semantic audio Transformer (HTS-AT) [2] for the text and audio encoders, respectively. We split the train data collected in Section 2 into 1,925 and 458 audio–text pairs for model training and validation.
We evaluated the correlation between CLAPScore and subjective evaluation scores using 2,405 audio–text pairs collected from the AudioCaps [12] test set in Section 2. We used SRCC and LCC to calculate the correlation between CLAPScore and subjective score, and used MSE to evaluate the error between CLAPScore and subjective score. LAION CLAP [26] and MS CLAP [5] are used as the comparison methods.
@@ 4.2 Results
Model Sub. score $\leq 5$ Sub. score >5>5